The following are two distinct demand functions: π‘ž1=24βˆ’0.2𝑃1 and π‘ž2=10βˆ’0.05𝑃2. The average cost AC=40+35/q. What price the firm will change without discrimination?


1
Expert's answer
2022-06-07T17:34:15-0400

"q_1=24-0.2p,"

"q_2=10-0.05p,"

"q_1=q_2,"

"24-0.2p=10-0.05p,"

"14=0.15p,"

"p=93.3,"

"q_1=q_2=5.3."